# Two Micron-Size Dark Dimensions

### v1

@@ Abstract
Two extra dimensions of micron scale might simultaneously address the gauge and cosmological hierarchy problems. In our paper we examine various observational bounds in scenarios with one and two large extra dimensions, to see if they are compatible with the micron scale. We show that consistency with astrophysical and cosmological observations requires that two extra dimensions of micron scale must not admit isometries, whereby conservation of the extra dimensional momentum is violated to a large extent, allowing thus the massive Kaluza-Klein modes of the graviton to promptly decay to other lighter graviton modes. We also show that within this scenario primordial black holes with masses in the range $10^{8}\lesssim M_{\rm BH}/{\rm g}\lesssim 10^{21}$ could make all cosmological dark matter.
@@ I General Idea
Extra dimensions are in general thought to be compact, since a four-dimensional (4D) description fits the world we perceive and measure (so far) very well. It is this compactness that sets a new length scale. A particularly attractive framework has as its main feature large extra dimensions along which only gravity propagates, because Standard Model (SM) fields are localized on a brane extended in the non-compact dimensions Arkani-Hamed:1998jmv . Within this framework the usual 4D graviton is complemented by a tower of Kaluza-Klein (KK) modes, corresponding to the new available phase space in the bulk. In the canonical example, spacetime is a direct product of a non-compact 4D spacetime manifold and a flat spatial $n$-torus of common linear size $2\pi R_{\rm KK}$ and volume $V_{n}=(2\pi R_{\rm KK})^{n}$. The $(4+n)$ dimensional Planck scale (or more generally the species scale $\Lambda_{\rm sp}$ where gravity becomes strong Dvali:2007hz ; Dvali:2007wp ; Dvali:2009ks ; Dvali:2010vm ; Dvali:2012uq ) is related to the 4D reduced Planck mass $M_{p}$ according to
@@
Most notably, this framework can be embedded into string theory Antoniadis:1998ig .
@@
selects $\alpha=1/4$ as the preferred value and gives in this way a procedure to connect $R_{\rm KK}$ to the dark energy scale $\Lambda/M_{p}^{4}\sim 10^{-120}$ (i.e. $\Lambda^{-1/4}\sim 2.39~{}{\rm meV}$ Planck:2018vyg ) as follows Montero:2022prj
@@
Actually, the relation (4) is independent of the number $n$ of dark dimensions of micron size. Taking into account the supernovae and neutron star constraints (see the discussion in the next section), the authors of Montero:2022prj have argued in favour of a single dark dimension in the micron range. In this case, one is led to a species scale of order Montero:2022prj
@@
The layout of the paper is as follows. In Sec. II we confront (6) to astrophysical data and in Sec. III to cosmological observations. In Sec. IV we reexamine whether neutrino masses and mixings could exclusively occur due to physics in the bulk. In Sec. V we show that an all-dark-matter interpretation in terms of 6D primordial black holes (PBH) should be feasible for masses in the range $10^{8}\lesssim M_{\rm BH}/{\rm g}\lesssim 10^{21}$. In Sec. VI we take a look at the possibility of probing $\Lambda_{\rm sp}$ with ultra-high-energy cosmic rays. Conclusions are given in Sec. VII.
II Astrophysical Constraints
@@
$\Gamma^{l}_{\gamma\gamma}=\frac{{\tilde{\lambda}}^{2}\ m^{3}_{\rm KK}\ l^{3}}{80\pi M_{p}^{2}}\,,$ (8)
where we have considered a tower of equally spaced gravitons, indexed by an integer $l$ and with mass $m_{l}=l\ m_{\rm KK}$, and where the parameter $\tilde{\lambda}$ measures the value of the dark graviton wave function at the SM brane and is expected to be ${\cal O}(1)$ Han:1998sg ; Hall:1999mk . The non-observation of these gamma rays in data collected by the Energetic Gamma Ray Experiment Telescope (EGRET) leads to the following limits on the compactification scale:
@@
    $R_{\rm KK}<\left\{\begin{array}[]{ll}4.9~{}{\rm m}&~{}~{}~{}~{}~{}~{}{\rm for}\ n=1\\ 96~{}{\rm nm}&~{}~{}~{}~{}~{}~{}{\rm for}\ n=2\end{array}\right.\,,$ (9)
@@
    $R_{\rm KK}<\left\{\begin{array}[]{ll}410~{}{\rm m}&~{}~{}~{}~{}~{}~{}{\rm for}\ n=1\\ 0.36~{}\mu{\rm m}&~{}~{}~{}~{}~{}~{}{\rm for}\ n=2\end{array}\right.\,,$ (10)
    corresponding to limits of $\Lambda_{\rm sp}>770~{}{\rm TeV}$ and $\Lambda_{\rm sp}>14~{}{\rm TeV}$, respectively Hannestad:2003yd .
@@
    $R_{\rm KK}<\left\{\begin{array}[]{ll}7.4~{}{\rm m}&~{}~{}~{}~{}~{}~{}{\rm for}\ n=1\\ 51~{}{\rm nm}&~{}~{}~{}~{}~{}~{}{\rm for}\ n=2\end{array}\right.\,,$ (11)
    corresponding to limits of $\Lambda_{\rm sp}>2.9\times 10^{3}~{}{\rm TeV}$ and $\Lambda_{\rm sp}>39~{}{\rm TeV}$, respectively Hannestad:2003yd .
@@
$R_{\rm KK}<\left\{\begin{array}[]{ll}44~{}\mu{\rm m}&~{}~{}~{}~{}~{}~{}{\rm for}\ n=1\\ 0.15~{}{\rm nm}&~{}~{}~{}~{}~{}~{}{\rm for}\ n=2\end{array}\right.\,,$ (12)
corresponding to $\Lambda_{\rm sp}>1.6\times 10^{5}~{}{\rm TeV}$ and $\Lambda_{\rm sp}>701~{}{\rm TeV}$, respectively Hannestad:2003yd .
Notwithstanding, the constraints from EGRET data and neutron-star heating rely on the assumption that KK modes can only decay into SM fields localized on the brane, but cannot decay into other KK modes with smaller bulk momenta. If large extra dimensions do not admit isometries this assumption is in general not valid and the bounds in (9), (10), (11), and (12) can be avoided Mohapatra:2003ah . To demonstrate this point we assume that the KK modes acquire a non-zero vacuum expectation value, $\langle\varphi_{\vec{n}}\rangle$, which is independent of $l$ and leads to a spontaneous breakdown of the translational invariance in the compact space. If this were the case, the violation of KK momentum conservation would allow intra-tower decays. The total decay width of $l$ KK graviton is found to be Mohapatra:2003ah
$\Gamma_{\rm tot}^{l}\sim\Gamma^{l}_{\gamma\gamma}\left(\frac{\langle\varphi_{\vec{n}}\rangle\ M_{p}}{\Lambda_{\rm sp}^{2}}\right)^{2}\,.$ (13)
Note that for $\langle\varphi_{\vec{n}}\rangle\sim\Lambda_{\rm sp}\sim 10~{}{\rm TeV}$, the total decay rate (13) is about 28 orders of magnitude larger than (8). As a consequence, the $l$ KK graviton would predominantly decay into lighter KK gravitons.
It was recently noted that the amount of energy that can be absorbed by the dark matter halo constrains the fraction of the supernova energy which could released via the form of dark radiation to roughly $10^{-5}$ Vogl:2024ack . However, we note that the efficiency of absorbing the free-streaming gravitons is extraordinarily small.
We conclude that in the absence of isometries the SN constraints are given by (7) and two micron size dimensions with $R_{\rm KK}\sim 1~{}\mu{\rm m}$ are possible. On the other hand, if the bounds (11) and (12) are applied, the only possible case is $n=1$ corresponding to one micron size dark dimension.
III Cosmological Constraints
Most of the cosmological constraints on large extra dimensions have also been derived on the assumption that the KK momentum of the graviton is conserved (see e.g. Hall:1999mk ; Fairbairn:2001ct ; Hannestad:2001nq ), and therefore these bounds can be avoided if the compact space is not translational invariant. We have seen that if the KK momentum of the graviton tower is not conserved a given KK mode of the tower could decay into final states that include other, lighter KK excitations. Such a “dark-to-dark” decay process usually dominates the cosmological evolution of the graviton tower and could be regarded as the telltale signature of the dynamical dark matter framework Dienes:2011ja .
A specific realization of dynamical dark matter has been proposed for $n=1$, assuming that a parent particle with mass $m_{l}$ can decay to two daughter particles with masses $m_{l^{\prime}}$ and $m_{l^{\prime\prime}}$ such that $m_{l}=m_{l^{\prime}}+m_{l^{\prime\prime}}+\epsilon$, with $\epsilon\leq m_{\rm KK}\delta$ and $\delta\sim{\cal O}(1)$ Gonzalo:2022jac . As might be expected, the KK mode $l$ decays with gravitational strength to lighter elements of the tower and so the partial decay width is shown to be
@@ III Cosmological Constraints
In principle, there are $(m_{l}/m_{\rm KK})^{2}$ possibilities for the $l^{\prime}l^{\prime\prime}$ pair of KK modes that $l$ can decay to, but because of a small violation of KK quantum number, the choice effectively becomes $m_{l}\delta/m_{\rm KK}$. Taking into account the phase factor which is roughly the velocity of decay products at threshold,
@@
$m_{l}\sim\left(\frac{M_{p}^{4}\ m_{\rm KK}}{\beta^{4}\ \delta^{3}}\right)^{1/7}t^{-2/7}\,,$ (17)

### v2

@@ Abstract
Two extra dimensions of micron scale might simultaneously address the gauge and cosmological hierarchy problems. In our paper we examine various observational bounds in scenarios with one and two large extra dimensions, to see if they are compatible with the micron scale. We show that consistency with astrophysical observations requires that two extra dimensions of micron scale must not admit isometries, whereby conservation of the extra dimensional momentum is violated, allowing the massive Kaluza-Klein modes of the graviton to decay to other lighter graviton modes. However, to remain consistent with cosmological observations two extra dimensions of micron scale require a delicately fine tuning of the temperature at which the universe enters the radiation dominated epoch. Diving into this fine-tuned scenario we also show that primordial black holes with masses in the range $10^{8}\lesssim M_{\rm BH}/{\rm g}\lesssim 10^{21}$ could make all cosmological dark matter.
@@ I General Idea
Extra dimensions are in general thought to be compact, since a four-dimensional (4D) description fits the world we perceive and measure (so far) very well. It is this compactness that sets a new length scale. A particularly attractive framework has as its main feature large extra dimensions along which only gravity propagates, because Standard Model (SM) fields are localized on a brane extended in the non-compact dimensions Arkani-Hamed:1998jmv . Most notably, this construct can be embedded into string theory Antoniadis:1998ig . Within this framework the usual 4D graviton is complemented by a tower of Kaluza-Klein (KK) modes, corresponding to the new available phase space in the bulk. In the canonical example, spacetime is a direct product of a non-compact 4D spacetime manifold and a flat spatial $n$-torus of common linear size $2\pi R_{\rm KK}$ and volume $V_{n}=(2\pi R_{\rm KK})^{n}$. The $(4+n)$ dimensional Planck scale (or more generally the species scale $\Lambda_{\rm sp}$ where gravity becomes strong Dvali:2007hz ; Dvali:2007wp ; Dvali:2009ks ; Dvali:2010vm ; Dvali:2012uq ) is related to the 4D reduced Planck mass $M_{p}$ according to
@@
selects $\alpha=1/4$ as the preferred value and gives in this way a procedure to connect $R_{\rm KK}$ to the dark energy scale $\Lambda/M_{p}^{4}\sim 10^{-120}$ (i.e. $\Lambda^{1/4}\sim 2.39~{}{\rm meV}$ Planck:2018vyg ) as follows Montero:2022prj
@@
Actually, the relation (4) is independent of the number $n$ of dark dimensions of micron size. Taking into account the supernovae and neutron star constraints (see the discussion in the next section), the authors of Montero:2022prj have argued in favor of a single dark dimension in the micron range. In this case, one is led to a species scale of order Montero:2022prj
@@
The layout of the paper is as follows. In Sec. II we confront (6) to null results from collider searches, in Sec. III we confront it to astrophysical data, and in Sec. IV we confront it to cosmological observations. In Sec. V we reexamine whether neutrino masses and mixings could exclusively occur due to physics in the bulk. In Sec. VI we show that an all-dark-matter interpretation in terms of 6D primordial black holes (PBH) should be feasible for masses in the range $10^{8}\lesssim M_{\rm BH}/{\rm g}\lesssim 10^{21}$. In Sec. VII we take a look at the possibility of probing $\Lambda_{\rm sp}$ with ultra-high-energy cosmic rays. Conclusions are given in Sec. VIII.
II Collider Constraints
ATLAS and CMS are the two primary, general-purpose experiments, at the Large Hadron Collider (LHC). Since the early days of the LHC the ATLAS and CMS collaborations developed analysis strategies to search for signals of large extra dimensions.
A particular signal of these searches is the production of a single very energetic “mono-object” that does not balance the transverse momentum carried by anything else emerging from the collision (as would be required by momentum and energy conservation). Examples of such objects are very energetic photons or particle jets. Collisions producing these mono-objects only appear to be imbalanced because the emerging photon or jet is balanced by a graviton that escapes detection. As a consequence, SM processes such as the production of a jet plus a $Z$ boson which decays into neutrinos can mimic a graviton production signal. The absence of any excess in the mono-photon or mono-jet event channels at ATLAS and CMS has lead to stringent limits on the species scale, see e.g. CMS:2012lmn ; ATLAS:2012ezx ; CMS:2011esc ; ATLAS:2011kno ; CMS:2014jvv ; ATLAS:2015qlt ; ATLAS:2021kxv ; CMS:2021far . Null results from searches of mono-objects were also reported by the Tevatron D0 D0:2008ayi and CDF CDF:2008njt collaborations.
One can also search for virtual graviton effects, which manifest themselves as a new contribution to the continuum in the invariant mass spectrum of two energetic photons or fermions (dileptons or dijets). As of today no signals have been observed at ATLAS and CMS, excluding such contributions for $\Lambda_{\rm sp}$ up to several TeV CMS:2018ucw ; ATLAS:2017ayi ; CMS:2018dqv ; CMS:2021ctt .
Today, the most restrictive bounds on the species scale come from searches in events with an energetic jet and missing transverse momentum. For $n=2$, the 95% CL lower bound reported by the ATLAS Collaboration is ATLAS:2021kxv
$\Lambda_{\rm sp,min}=4.5~{}{\rm TeV}$ (7)
whereas the CMS Collaboration reported CMS:2021far
$\Lambda_{\rm sp,min}=4.5~{}{\rm TeV}\,.$ (8)
We note that we have reduced the limits as given in in Table 10 of ATLAS:2021kxv and in Table 3 of CMS:2021far by a factor of $(2\pi)^{n/(n+2)}$. This is because both the ATLAS ATLAS:2015qlt and CMS CMS:2014jvv collaborations definine the volume of the compact space as introduced in Giudice:1998ck , i.e., volume equal to $R_{\rm KK}^{n}$ rather than $(2\pi R_{\rm KK})^{n}$.
If the large extra dimension scenario is embedded in a string theory at the TeV scale Antoniadis:1998ig , we expect the string scale $M_{s}\lesssim\Lambda_{\rm sp}$, and thus production of string resonances at the LHC Anchordoqui:2007da ; Anchordoqui:2008ac ; Lust:2008qc ; Anchordoqui:2008di ; Anchordoqui:2009mm ; Anchordoqui:2014wha . Only one assumption is necessary in order to set up a solid framework: the string coupling must be small in order to rely on perturbation theory in the computations of scattering amplitudes. In this case, black hole production and other strong gravity effects occur at energies above the string scale; therefore at least a few lowest Regge recurrences are available for examination, free from interference with some complex quantum gravitational phenomena. Certain amplitudes to leading order in string coupling (but including all string $\alpha^{\prime}$ corrections) are universal. These amplitudes, which include $2\to 2$ scattering processes involving 4 gluons or 2 gluons and 2 quarks, are independent of the details of the compactification, such as the configuration of branes, the geometry of the extra dimensions, and whether SUSY is broken or not. This model independence makes it possible to compute the string corrections to dijet signals at the LHC Anchordoqui:2008di .
In addition, the mixing between hypercharge and the gauge baryon number symmetry implies that tree level gluon-gluon scattering processes such as $gg\to g\gamma$ and $gg\to\gamma\gamma$ are allowed even though they can only appear at loop order in the SM Anchordoqui:2007da . Unlike the dijet case, these amplitudes are mildly model dependent, i.e., they depend on a mixing parameter which is fixed by the D-brane model but are otherwise independent of the compactification scheme.
The ATLAS and CMS collaborations have searched for signals of Regge recurrences analyzing the dijet and $\gamma$ + jet invariant mass distributions Khachatryan:2010jd ; Chatrchyan:2011ns ; ATLAS:2011ai ; Chatrchyan:2013qha ; Aad:2013cva ; Khachatryan:2015sja ; Khachatryan:2015dcf ; Sirunyan:2016iap ; Sirunyan:2018xlo ; Sirunyan:2019vgj . Null results of these searches lead to a lower bound $M_{s}>8~{}{\rm TeV}$ at the 95% CL. Putting all this together, we conclude that collider constraints are consistent with two dark dimensions of micron size and $\Lambda_{\rm sp}\sim 10~{}{\rm TeV}$.
III Astrophysical Constraints
@@
[…]
where we have considered a tower of equally spaced gravitons, indexed by an integer $\vec{l}$ and with mass $m_{l}=\vec{l}\ m_{\rm KK}$, and where the parameter $\tilde{\lambda}$ measures the value of the dark graviton wave function at the SM brane and is expected to be ${\cal O}(1)$ Han:1998sg ; Hall:1999mk . Following Hannestad:2003yd we assume $\tilde{\lambda}=1$. The non-observation of these gamma rays in data collected by the Energetic Gamma Ray Experiment Telescope (EGRET) leads to the following limits on the compactification scale:
@@
    $R_{\rm KK}<\left\{\begin{array}[]{ll}4.9\times 10^{2}~{}{\rm m}&~{}~{}~{}~{}~{}~{}{\rm for}\ n=1\\ 1~{}\mu{\rm m}&~{}~{}~{}~{}~{}~{}{\rm for}\ n=2\end{array}\right.\,,$ (11)
@@
    $R_{\rm KK}<\left\{\begin{array}[]{ll}5.3\times 10^{3}~{}{\rm m}&~{}~{}~{}~{}~{}~{}{\rm for}\ n=1\\ 1.3~{}\mu{\rm m}&~{}~{}~{}~{}~{}~{}{\rm for}\ n=2\end{array}\right.\,,$ (12)
    corresponding to limits of $\Lambda_{\rm sp}>330~{}{\rm TeV}$ and $\Lambda_{\rm sp}>7.7~{}{\rm TeV}$, respectively Hannestad:2003yd .
@@
    $R_{\rm KK}<\left\{\begin{array}[]{ll}16~{}{\rm m}&~{}~{}~{}~{}~{}~{}{\rm for}\ n=1\\ 67~{}{\rm nm}&~{}~{}~{}~{}~{}~{}{\rm for}\ n=2\end{array}\right.\,,$ (13)
    corresponding to limits of $\Lambda_{\rm sp}>2.2\times 10^{3}~{}{\rm TeV}$ and $\Lambda_{\rm sp}>34~{}{\rm TeV}$, respectively Hannestad:2003yd .
@@
$R_{\rm KK}<\left\{\begin{array}[]{ll}8.3~{}{\rm m}&~{}~{}~{}~{}~{}~{}{\rm for}\ n=1\\ 59~{}{\rm nm}&~{}~{}~{}~{}~{}~{}{\rm for}\ n=2\end{array}\right.\,,$ (14)
corresponding to $\Lambda_{\rm sp}>2.8\times 10^{3}~{}{\rm TeV}$ and $\Lambda_{\rm sp}>36~{}{\rm TeV}$, respectively Hannestad:2003yd .
Notwithstanding, the constraints coming from the decays of KK gravitons into photons are model dependent and can be completely evaded. Note that the constraints from EGRET data and neutron-star heating rely on the assumption that KK modes can only decay into SM fields localized on the brane, but cannot decay into other KK modes with smaller bulk momenta. If large extra dimensions do not admit isometries this assumption is in general not valid and the bounds in (11), (12), (13), and (14) can be evaded Mohapatra:2003ah . Indeed, if the KK momentum of the graviton tower is not conserved a given KK mode of the tower could decay into final states that include other, lighter KK excitations.
Following Gonzalo:2022jac , we set $n=1$ and assume that a parent particle with mass $m_{l}$ can decay to two daughter particles with masses $m_{l^{\prime}}$ and $m_{l^{\prime\prime}}$ such that $m_{l}=m_{l^{\prime}}+m_{l^{\prime\prime}}+\epsilon$, with $\epsilon\leq m_{\rm KK}\delta$ and $\delta\sim{\cal O}(1)$. As might be expected, the KK mode $l$ decays with gravitational strength to lighter elements of the tower and so the partial decay width is shown to be
@@
In principle, there are $m_{l}/m_{\rm KK}$ possibilities for the $l^{\prime}l^{\prime\prime}$ pair of KK modes that $l$ can decay to, but because of a small violation of KK quantum number, the choice effectively becomes $m_{l}\delta/m_{\rm KK}$. Taking into account the phase space factor which is roughly the velocity of decay products at threshold,
@@
$m_{l}(t)\sim\left(\frac{M_{p}^{4}\ m_{\rm KK}}{\beta^{4}\ \delta^{3}}\right)^{1/7}t^{-2/7}\,,$ (18)
@@
For $n=2$, conservation of energy forces the the momenta of the decay products to be almost parallel and so the number of decay channels available is still effectively 1D, up to a width that will be set by $\delta$. The evolution of the KK mass in the graviton tower is also given by (18).
We now turn to discuss how dark-to-dark decays provide a path to evade the lower bounds on the species scale given in (11), (13), and (14). Following Hannestad and Raffelt (HR), we parametrize the bound on the species scale in the absence of intra-tower decays by
[…]
$\Lambda_{\rm sp,min}^{\rm HR}=\left(\frac{f_{\rm KK}^{\rm HR}}{0.5}\right)^{-1/4}8.9~{}{\rm TeV}\,.$ (20)
Due to the intra-tower decays the fraction of KK gravitons that could decay into photons gets a multiplicative correction to accomodate the evolution of dark-to-dark decays,
$f_{\rm KK}^{\rm new}=f_{\rm KK}^{\rm HR}\times\overline{f_{\rm KK}}$ (21)
where $\overline{f_{\rm KK}}$ is determined by the ratio of the initial decay rate $\Gamma_{0}$ of the KK gravitons to photons to the final decay rate $\Gamma(t)$ of the KK gravitons to photons,
$\overline{f_{\rm KK}}=\int_{0}^{\rm\tau_{\rm NS}}\frac{\Gamma_{0}}{\Gamma(t)}\ \frac{dt}{\tau_{\rm NS}}\,,$ (22)
and where $\tau_{\rm NS}\sim 17~{}{\rm Myr}$ is the lifetime of the neutron star Hannestad:2003yd ,
$\Gamma_{0}\sim m_{0}^{3}/M_{p}^{2}\,,$ (23)
$m_{0}\sim 100~{}{\rm MeV}$ is the mass of the KK modes at production, and
$\Gamma(t)=m_{l}^{3}(t)/M_{p}^{2}\,.$ (24)
Substituting (23) and (24) into (22) while using (18) we obtain
$\overline{f_{\rm KK}}\sim\frac{13}{7}\left[\frac{m_{0}}{(M_{p}^{4}\ m_{\rm KK})^{1/7}}\right]^{3}\tau_{\rm NS}^{6/7}\sim 881\,,$ (25)
where we have taken $\delta=\beta=1$. The corrected lower limit on the species scale is then
[…]
All in all, for $n=2$, the new bounds on the species scale associated to (13) and (14) become
[…]
and
[…]
respectively.
Duplicating this procedure, but substituting $\tau_{\rm NS}$ by the age of the universe in the limit of integration of (22), we obtain the $\overline{f_{\rm KK}}$ to rescale the bound on the species scale associated to (11). It is straightforward to see that the corresponding value of $\overline{f_{\rm KK}}>881$ and therefore the bound on (11) can be evaded.
We conclude that in the absence of isometries the SN constraints are given by (9) and one or two extra dimensions with $R_{\rm KK}\sim 1~{}\mu{\rm m}$ are consistent with observations.
IV Cosmological Constraints
Cosmological observations lead to additional model independent and model dependent constraints. We begin discussing constraints on models in which KK intra-tower decay is not allowed, and then we relax this condition by considering extra dimensions which do not admit isometries.
Model independent constraints originate again in the emission of KK modes. Expansion dominated cooling gives a model independent bound. Indeed, an upper limit on the “normalcy” temperature $T_{*}$ at which the universe must be free of bulk modes can be derived by demanding that the rate at which radiation energy density on the brane evaporates into KK modes in the bulk remains below the normal cooling due to cosmological expansion,
$T_{*}\lesssim 10^{\frac{7n-7}{n+1}}\ \left(\frac{\Lambda_{\rm sp}}{10~{}{\rm TeV}}\right)^{\frac{n+2}{n+1}}~{}{\rm MeV}\,,$ (29)
which for $n=2$, leads to $T_{*}\lesssim 200~{}{\rm MeV}$ Arkani-Hamed:1998sfv . Big Bang nucleosynthesis (BBN) set a more restrictive (but model dependent) bound on $T_{*}$. From the 4D perspective, the gravitons produced at temperature $T$ are massive KK modes with mass $\sim T$. This is because the emission rate of each KK mode $\sim T^{3}/M_{p}$. It goes without saying that for given $T$, all KK modes up to $T$ are produced, but most particles have a mass $\sim T$, because the emission rate goes down rapidly with decreasing $T$. Putting all this together, the energy density in KK modes redshifts away as $a^{3}$ rather than $a^{4}$, where $a$ is the cosmic scale factor. Then, to insure normal expansion rate during BBN the bound on the normalcy temperature is slightly stronger,
$T_{*}\lesssim 10^{\frac{7n-7}{n+2}}\ \left(\frac{\Lambda_{\rm sp}}{10~{}{\rm TeV}}\right)~{}{\rm MeV}\,,$ (30)
which for $n=2$, leads to $T_{*}\lesssim 50~{}{\rm MeV}$ Arkani-Hamed:1998sfv .
Even more restrictive bounds emerge by requiring that the energy density of KK gravitons today does not overclose the universe. Indeed, to overcome overproduction of KK modes at early times the temperature at which the universe enters the radiation dominated epoch must be delicately fine-tuned Arkani-Hamed:1998sfv ; Hall:1999mk ; Hannestad:2001nq . The contribution of thermally produced KK to the energy density today is estimated to be
$\frac{\rho_{{\rm KK}_{0}}}{T_{0}^{3}}\sim\frac{2}{3}\frac{M_{p}T_{*}^{4}}{\sqrt{g_{*}}\Lambda_{\rm sp}^{4}}\,,$ (31)
where $g_{*}$ is the effective number of relativistic degrees of freedom and where we have used the fact that the ratio $\rho_{\rm KK}/T^{3}$ is invariant. The critical density of the universe today corresponds to $(\rho_{\rm crit}/T_{0}^{3})\sim 3\times 10^{-9}~{}{\rm GeV}$. Setting $\rho_{\rm KK}\leq\rho_{\rm crit}$, with $\Lambda_{\rm sp}=10~{}{\rm TeV}$ and $g_{*}=10.75$ leads $T_{*}\sim 2.8~{}{\rm MeV}$ Arkani-Hamed:1998sfv . This normalcy temperature derived more carefully is closer to the decoupling temperature for the muon and tau neutrinos. Indeed, for $T_{*}=2.15~{}{\rm MeV}$, it follows that
$R_{\rm KK}<3.3\ h~{}\mu{\rm m}\sim 2.3~{}\mu{\rm m}\,,$ (32)
which corresponds to
$\Lambda_{\rm sp}>5~{}{\rm TeV}/\sqrt{h}\sim 6~{}{\rm TeV}\,,$ (33)
where $H_{0}=100\,h~{}{\rm km\,Mpc^{-1}\,s^{-1}}$ is the Hubble constant, with $h\simeq 0.7$ Hall:1999mk . [^1] Note that $M=2^{1/(n+2)}(2\pi)^{n/(n+2)}\Lambda_{\rm sp}$ Hannestad:2003yd , where $M$ is the underlying $(n+4)$-dimensional scale adopted in Hall:1999mk . Needless to say, the bounds in (32) and (33) only take into account thermal production of the KK modes of gravitons in the early universe. Successful inflation and reheating, as well as baryogenesis, usually stand in need for the existence of fields in the bulk, most notably the inflaton. The non-thermal production of KK modes accompanying the inflaton decay would further constrain $R_{\rm KK}$ and exclude the possibility of two extra dimensions of micron size Hannestad:2001nq . [^2] For values of $T_{*}$ above the QCD phase transition further considerations are necessary to constrain $R_{\rm KK}$ Fairbairn:2001ct ; Macesanu:2004gf .
If we are prepared to accept that KK gravitons are only produced thermally, it is of interest to further investigate whether low reheating temperature universes, with $T_{*}\sim 2~{}{\rm MeV}$, remain consistent with cosmological observations. Lower bounds on the reheating temperature have been derived assuming the the late-time entropy production near BBN is dominated by radiative decays. The incomplete thermalization of neutrinos could modify the production of 4He and this leads to a bound $T_{*}>0.5~{}{\rm MeV}$ at 95%CL Kawasaki:1999na . When hadronic decays are also considered $T_{*}>2.4~{}{\rm MeV}$ at 95%CL Kawasaki:2000en . In addition, the effect of flavor neutrino oscillations turns out to be quite relevant to accommodate BBN, shifting the lower bound to $T_{*}$ higher temperatures. Indeed, the thermalization process of two- and three-flavor neutrino oscillations have been analyzed yielding the following 95% CL bounds: $T_{*}>2~{}{\rm MeV}$ and $T_{*}>4.1~{}{\rm MeV}$ assuming radiative decay of the massive particles Ichikawa:2005vw ; deSalas:2015glj . More recently, it was shown that when both neutrino oscillations and self-interactions are considered in the thermalization process the 95% CL lower bounds on the normalcy temperature are less restrictive: $T_{*}>1.8~{}{\rm MeV}$ and $T_{*}>4~{}{\rm MeV}$, for radiative decays and hadronic decays; respectively Hasegawa:2019jsa . CMB data and large scale structure lead to comparable bounds on the normalcy temperature, $T_{*}\gtrsim 2~{}{\rm MeV}$ Hannestad:2004px ; Ichikawa:2006vm ; DeBernardis:2008zz . [^3] We note in passing that a recent study combining Planck CMB data and DESI DR1 results with $N_{\rm eff}=2.58$ gives $T_{*}>3.79~{}{\rm MeV}$ at 95% CL, and when BBN predictions are included in the likelihood analysis using $N_{\rm eff}=2.98$ yields $T_{*}>5.96~{}{\rm MeV}$ Barbieri:2025moq . However, it should also be noted that because both the primordial helium abundance $Y_{\rm P}$ and the relativistic degrees of freedom affect the CMB damping tail, they are partially degenerate. Allowing $N_{\rm eff}$ and $Y_{\rm P}$ to vary in the likelihood analysis leads to weaker lower bound of $N_{\rm eff}=2.32$ Planck:2018vyg . This will naturally shift the value of the minimum allowed $T_{*}$ to lower temperatures.
Recently, cosmologies with ${\cal O}({\rm MeV})$ normalcy temperatures have become the focus of a great deal of interest because they can accommodate several new physics scenarios that would normally be constrained by high-temperature reheating models, including massive sterile neutrinos Abazajian:2023reo . For example, if $T_{*}\sim 1.8~{}{\rm MeV}$ sterile neutrinos (of mass $m_{s}\sim{\rm keV}$) with large mixing angle $\theta$ and a dark decay to radiation may ameliorate the $H_{0}$ tension Escudero:2022rbq . Several neutrino experiments are scanning the 0.1 to 100 keV mass range to search for these sterile neutrinos, which may be visible in the lab but may be invisible to cosmological observations; for details see Fig. 1.
Figure 1: The $(m_{s},\theta)$ parameter space for $T_{*}=1.8~{}{\rm MeV}$. The black region shows constraints from a collection of nuclear beta decay searches ParticleDataGroup:2016lqr and the region in golden yellow constraints from $\beta$-decay searches at BeEST Friedrich:2020nze . The thermal nature of the CMB constrains the red hatched portion in the upper right Hu:1993gc . In the pink region $1\leq N_{\rm eff}\leq 3.044$. The diagonal red line correspond to the scenario when the decay happens at the temperature of matter-radiation equality. Above the diagonal red line, $N_{\rm eff}>3.044$ and the parameter space can accommodate an alleviation of the $H_{0}$ tension. The experimental sensitivity of HUNTER Smith:2016vku , TRISTAN KATRIN:2018oow , MAGNETO-$\nu$ Lee:2025txi , and PTOLEMY PTOLEMY:2019hkd for nuclear decay searches of keV-scale sterile neutrinos is also shown. Taken from Abazajian:2023reo .
For $n$ extra dimensions, the constraint from cosmological overproduction of bulk graviton modes can be written as
$T_{*}\sim 10^{3+(6n-15)/(n+2)}\left(\frac{\Lambda_{\rm sp}}{10^{9}~{}{\rm GeV}}\right)~{}{\rm GeV}\,,$ (34)
which implies that for a single dark dimension in the micron range the universe must be free of bulk modes at $T_{*}\sim 1~{}{\rm GeV}$ Arkani-Hamed:1998sfv .